OpenStudy (anonymous):

Given the function f(x) = 2(x + 10), find x if f(x) = 24.

OpenStudy (anonymous):

Is the answer 7?

Parth (parthkohli):

We know that \(f(x) = 24\). We also know that \(f(x) = 2(x + 10)\) So, \(2(x + 10) = 24\)

Parth (parthkohli):

And nope... not 7.

OpenStudy (anonymous):

after you distribute the 2. do you subtract 10 from 24 or 24 from 10?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

oh hold on i distributed it wrong.

Parth (parthkohli):

Don't distribute—divide both sides by 2.\[2(x + 10) = 24\implies {2(x + 10) \over 2} = 24 \]Okay?

Parth (parthkohli):

Oops...\[{2(x + 10) \over 2} = {24 \over 2} \]

OpenStudy (anonymous):

ohh i see! i got 2 as my answer but i still distributed

Parth (parthkohli):

Works both ways. :3

OpenStudy (anonymous):

oh so its 12?

Parth (parthkohli):

No. The answer still stays 2.

OpenStudy (anonymous):

okay. thank you!!

Parth (parthkohli):

You're welcome! c: